# ROSE: Register Assisted General Time Series Forecasting with Decomposed Frequency Learning

### v1

@@ Appendix D Model generality
Autoformer +Multi Freq Mask 0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
@@
TimesNet +Multi Freq Mask 0.4 0.406 0.291 0.333 0.458 0.45 0.414 0.427
@@
PatchTST +Multi Freq Mask 0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381
@@
Appendix E Visualization of forecasting results
To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 10 to Figures 13, given by the following models: DLinear [40], TimesNet [18], iTransfomrer [20], PatchTST [6], and TiDE [39]. Among the methods, ROSE demonstrates the most accurate prediction ability.
Figure 10: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset with full-sample.
Figure 11: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset with 10% samples.
Figure 12: Visualization of input-512 and predict-336 forecasting results on the weather dataset with full-sample.
Figure 13: Visualization of input-512 and predict-336 forecasting results on the weather dataset with 10%-sample.
Appendix F Limitations
Our pre-training datasets still have room for expansion. While ROSE demonstrates strong performance, its effectiveness could be enhanced with a larger and more diverse set of pre-training datasets. By collecting additional real-world scenario data for pre-training, we aim to further improve the robustness and generalizability of ROSE across various domains.
Appendix G Broader Impacts
This paper presents ROSE, a novel general time series forecasting model that demonstrates promising performance. ROSE performs well in few-shot scenarios and shows notable zero-shot capabilities, highlighting its scalability and efficiency. We believe it can be a valuable addition to the pre-training research community. To support future research, we will also release the codebase for time-series pre-training.
This paper only focuses on the algorithm design. Using all the codes and datasets strictly follows the corresponding licenses A.1. There is no potential ethical risk or negative social impact.
Appendix H Full Results
H.1 Full-shot results
@@
Models ROSE Timer ITransformer PatchTST FITS Timesnet TiDE Dlinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.275 0.328 0.339 0.383 0.334 0.368 0.293 0.346 0.303 0.352 0.338 0.375 0.306 0.349 0.299 0.343
192 0.324 0.358 0.396 0.430 0.377 0.391 0.333 0.370 0.337 0.368 0.374 0.387 0.335 0.366 0.335 0.365
336 0.354 0.377 0.451 0.471 0.426 0.420 0.369 0.392 0.366 0.385 0.410 0.411 0.364 0.384 0.369 0.386
720 0.411 0.407 0.558 0.540 0.491 0.459 0.416 0.420 0.415 0.411 0.478 0.450 0.413 0.413 0.425 0.421
avg 0.341 0.367 0.436 0.456 0.407 0.410 0.353 0.382 0.355 0.379 0.400 0.406 0.355 0.378 0.357 0.379
ETTm2 96 0.157 0.243 0.184 0.263 0.180 0.264 0.166 0.256 0.162 0.253 0.187 0.267 0.161 0.251 0.167 0.260
192 0.213 0.283 0.278 0.334 0.250 0.309 0.223 0.296 0.216 0.291 0.249 0.309 0.215 0.289 0.224 0.303
336 0.266 0.319 0.334 0.375 0.311 0.348 0.274 0.329 0.268 0.326 0.321 0.351 0.267 0.326 0.281 0.342
720 0.347 0.373 0.435 0.431 0.412 0.407 0.362 0.385 0.348 0.378 0.408 0.403 0.352 0.383 0.397 0.421
avg 0.246 0.305 0.308 0.351 0.288 0.332 0.256 0.317 0.249 0.312 0.291 0.333 0.249 0.312 0.267 0.332
ETTh1 96 0.354 0.385 0.358 0.345 0.386 0.405 0.370 0.400 0.372 0.402 0.384 0.402 0.375 0.398 0.375 0.399
192 0.389 0.407 0.438 0.448 0.441 0.436 0.413 0.429 0.404 0.422 0.436 0.429 0.412 0.422 0.405 0.416
336 0.406 0.422 0.491 0.484 0.487 0.458 0.422 0.440 0.427 0.438 0.491 0.469 0.435 0.433 0.439 0.443
720 0.413 0.443 0.644 0.559 0.503 0.491 0.447 0.468 0.424 0.454 0.521 0.500 0.454 0.465 0.472 0.490
avg 0.391 0.414 0.483 0.459 0.454 0.447 0.413 0.434 0.407 0.429 0.458 0.450 0.419 0.430 0.423 0.437
ETTh2 96 0.265 0.320 0.288 0.346 0.297 0.349 0.274 0.337 0.271 0.336 0.340 0.374 0.270 0.336 0.289 0.353
192 0.328 0.369 0.380 0.409 0.380 0.400 0.341 0.382 0.331 0.374 0.402 0.414 0.332 0.380 0.383 0.418
336 0.353 0.391 0.372 0.413 0.428 0.432 0.329 0.384 0.354 0.395 0.452 0.452 0.360 0.407 0.448 0.465
720 0.376 0.417 0.425 0.456 0.427 0.445 0.379 0.422 0.377 0.423 0.462 0.468 0.419 0.451 0.605 0.551
avg 0.330 0.374 0.366 0.406 0.383 0.407 0.331 0.381 0.333 0.382 0.414 0.427 0.345 0.394 0.431 0.447
Traffic 96 0.354 0.252 0.351 0.255 0.395 0.268 0.360 0.249 0.385 0.271 0.593 0.321 0.336 0.253 0.410 0.282
192 0.377 0.257 0.387 0.269 0.417 0.276 0.379 0.256 0.397 0.275 0.617 0.336 0.346 0.257 0.423 0.287
336 0.396 0.262 0.426 0.288 0.433 0.283 0.392 0.264 0.410 0.280 0.629 0.336 0.355 0.260 0.436 0.296
720 0.434 0.283 0.489 0.323 0.467 0.302 0.432 0.286 0.448 0.301 0.640 0.350 0.386 0.273 0.466 0.315
avg 0.390 0.264 0.413 0.284 0.428 0.282 0.391 0.264 0.410 0.282 0.620 0.336 0.356 0.261 0.434 0.295
Weather 96 0.145 0.182 0.154 0.208 0.174 0.214 0.149 0.198 0.143 0.195 0.172 0.220 0.166 0.222 0.176 0.237
192 0.183 0.226 0.243 0.287 0.221 0.254 0.194 0.241 0.186 0.238 0.219 0.261 0.209 0.263 0.220 0.282
336 0.232 0.267 0.329 0.345 0.278 0.296 0.245 0.282 0.236 0.277 0.280 0.306 0.254 0.301 0.265 0.319
720 0.309 0.327 0.384 0.384 0.358 0.347 0.314 0.334 0.307 0.330 0.365 0.359 0.313 0.340 0.323 0.362
avg 0.217 0.251 0.277 0.306 0.258 0.278 0.226 0.264 0.218 0.260 0.259 0.287 0.236 0.282 0.246 0.300
Solar 96 0.171 0.199 0.230 0.245 0.203 0.237 0.190 0.273 0.198 0.253 0.250 0.292 0.200 0.270 0.206 0.281
192 0.185 0.215 0.249 0.261 0.233 0.261 0.204 0.302 0.218 0.262 0.296 0.318 0.219 0.285 0.225 0.291
336 0.193 0.218 0.255 0.265 0.248 0.273 0.212 0.293 0.235 0.270 0.319 0.330 0.234 0.296 0.240 0.300
720 0.198 0.224 0.272 0.276 0.249 0.275 0.221 0.310 0.241 0.273 0.338 0.337 0.242 0.303 0.248 0.307
avg 0.187 0.214 0.251 0.262 0.233 0.262 0.207 0.294 0.223 0.264 0.301 0.319 0.223 0.288 0.230 0.295
Electricity 96 0.125 0.220 0.136 0.230 0.148 0.240 0.129 0.222 0.134 0.231 0.168 0.272 0.132 0.229 0.140 0.237
192 0.142 0.235 0.168 0.260 0.162 0.253 0.147 0.240 0.149 0.244 0.184 0.289 0.147 0.243 0.153 0.249
336 0.162 0.252 0.192 0.285 0.178 0.269 0.163 0.259 0.165 0.260 0.198 0.300 0.161 0.261 0.169 0.267
720 0.191 0.284 0.230 0.322 0.225 0.317 0.197 0.290 0.203 0.293 0.220 0.320 0.196 0.294 0.203 0.301
avg 0.155 0.248 0.181 0.274 0.178 0.270 0.159 0.253 0.163 0.257 0.193 0.295 0.159 0.257 0.166 0.264
H.2 Few-shot results
@@
Models ROSE Timer ITransformer PatchTST FITS Timesnet TiDE Dlinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.287 0.336 0.351 0.398 0.353 0.392 0.317 0.363 0.311 0.354 0.481 0.446 0.373 0.388 0.454 0.475
192 0.331 0.362 0.429 0.448 0.385 0.410 0.351 0.382 0.351 0.378 0.621 0.491 0.406 0.403 0.575 0.548
336 0.362 0.379 0.496 0.495 0.422 0.432 0.376 0.398 0.377 0.396 0.521 0.479 0.436 0.425 0.773 0.631
720 0.416 0.412 0.573 0.573 0.494 0.472 0.435 0.430 0.441 0.436 0.571 0.508 0.499 0.457 0.943 0.716
avg 0.349 0.372 0.462 0.478 0.414 0.426 0.370 0.393 0.370 0.391 0.549 0.481 0.429 0.418 0.686 0.593
ETTm2 96 0.159 0.247 0.196 0.273 0.183 0.279 0.170 0.259 0.166 0.258 0.212 0.292 0.195 0.280 0.493 0.476
192 0.217 0.287 0.290 0.338 0.247 0.320 0.226 0.297 0.222 0.295 0.297 0.353 0.257 0.317 0.923 0.658
336 0.268 0.322 0.350 0.386 0.3 0.353 0.284 0.333 0.273 0.328 0.328 0.364 0.320 0.354 1.407 0.822
720 0.355 0.377 0.436 0.435 0.385 0.408 0.363 0.382 0.370 0.385 0.456 0.440 0.421 0.410 1.626 0.905
avg 0.249 0.308 0.318 0.358 0.279 0.340 0.261 0.318 0.258 0.317 0.323 0.362 0.298 0.340 1.112 0.715
ETTh1 96 0.367 0.395 0.361 0.409 0.442 0.464 0.458 0.463 0.438 0.449 0.579 0.522 0.708 0.569 1.355 0.816
192 0.399 0.416 0.398 0.427 0.476 0.475 0.481 0.490 0.512 0.489 0.641 0.553 0.744 0.584 1.210 0.825
336 0.405 0.423 0.405 0.436 0.486 0.482 0.465 0.475 0.584 0.532 0.721 0.582 0.768 0.599 1.487 0.914
720 0.416 0.443 0.492 0.503 0.509 0.506 0.478 0.492 0.681 0.590 0.630 0.574 0.770 0.606 1.369 0.826
avg 0.397 0.419 0.414 0.444 0.478 0.482 0.470 0.480 0.554 0.515 0.643 0.558 0.748 0.589 1.355 0.845
ETTh2 96 0.273 0.332 0.298 0.356 0.333 0.385 0.350 0.389 0.297 0.360 0.378 0.413 0.434 0.432 1.628 0.724
192 0.334 0.376 0.374 0.404 0.402 0.428 0.416 0.426 0.363 0.399 0.463 0.460 0.528 0.480 1.388 0.713
336 0.358 0.397 0.385 0.425 0.438 0.452 0.401 0.429 0.388 0.424 0.507 0.495 0.473 0.457 1.595 0.772
720 0.376 0.417 0.468 0.487 0.466 0.477 0.436 0.457 0.423 0.456 0.516 0.501 0.567 0.513 1.664 0.857
avg 0.335 0.380 0.381 0.418 0.410 0.436 0.401 0.425 0.368 0.410 0.466 0.467 0.501 0.471 1.569 0.766
Traffic 96 0.398 0.270 0.353 0.277 0.458 0.314 0.421 0.299 0.406 0.301 0.705 0.386 0.681 0.412 0.616 0.385
192 0.405 0.270 0.405 0.288 0.473 0.319 0.439 0.313 0.523 0.424 0.710 0.393 0.624 0.385 0.710 0.480
336 0.417 0.277 0.428 0.300 0.491 0.329 0.448 0.318 0.624 0.479 0.863 0.456 0.613 0.368 0.723 0.481
720 0.452 0.294 0.487 0.332 0.536 0.361 0.478 0.320 0.937 0.629 0.928 0.485 0.653 0.387 0.673 0.436
avg 0.418 0.278 0.418 0.299 0.490 0.331 0.447 0.312 0.623 0.458 0.801 0.430 0.643 0.388 0.680 0.446
Weather 96 0.145 0.184 0.180 0.228 0.189 0.229 0.166 0.217 0.174 0.222 0.199 0.248 0.208 0.240 0.230 0.318
192 0.190 0.227 0.256 0.296 0.239 0.269 0.211 0.257 0.226 0.265 0.249 0.285 0.249 0.274 0.357 0.425
336 0.245 0.269 0.321 0.344 0.294 0.308 0.261 0.296 0.279 0.301 0.297 0.316 0.283 0.306 0.464 0.493
720 0.317 0.328 0.420 0.398 0.366 0.356 0.328 0.342 0.361 0.354 0.367 0.361 0.365 0.354 0.515 0.532
avg 0.224 0.252 0.294 0.317 0.272 0.291 0.242 0.278 0.260 0.285 0.278 0.303 0.276 0.293 0.391 0.442
Solar 96 0.177 0.203 0.230 0.284 0.292 0.310 0.202 0.245 0.242 0.306 0.306 0.323 0.359 0.382 0.748 0.757
192 0.185 0.214 0.246 0.300 0.313 0.321 0.222 0.256 0.262 0.315 0.356 0.354 0.395 0.400 0.752 0.759
336 0.193 0.218 0.278 0.319 0.332 0.335 0.252 0.272 0.276 0.321 0.387 0.362 0.429 0.410 0.754 0.760
720 0.208 0.225 0.255 0.300 0.355 0.338 0.264 0.273 0.280 0.329 0.396 0.361 0.410 0.378 0.755 0.761
avg 0.191 0.215 0.252 0.301 0.323 0.326 0.235 0.262 0.265 0.318 0.361 0.350 0.398 0.392 0.752 0.759
Electricity 96 0.135 0.226 0.156 0.254 0.184 0.276 0.161 0.256 0.144 0.242 0.279 0.359 0.271 0.359 0.227 0.334
192 0.150 0.240 0.254 0.275 0.192 0.284 0.163 0.257 0.174 0.283 0.282 0.363 0.216 0.308 0.265 0.366
336 0.166 0.258 0.212 0.306 0.216 0.308 0.173 0.266 0.225 0.341 0.289 0.367 0.237 0.324 0.339 0.417
720 0.205 0.290 0.271 0.352 0.265 0.347 0.221 0.313 0.408 0.495 0.333 0.399 0.258 0.330 0.482 0.478
avg 0.164 0.253 0.223 0.297 0.214 0.304 0.180 0.273 0.238 0.341 0.296 0.372 0.245 0.330 0.328 0.399
H.3 Ablation study results

### v2

@@ A.7 Model generality
Autoformer +Multi-frequency Masking 0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
@@
TimesNet +Multi-frequency Masking 0.400 0.406 0.291 0.333 0.458 0.450 0.414 0.427
@@
PatchTST +Multi-frequency Masking 0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381
@@
A.8 Results deviation
We have conducted ROSE three times with different random seeds and have recorded the standard deviations for both the full-shot setting and the 10% few-shot setting, as illustrated in Table 11. As the baselines didn’t report deviations in the original paper, we only reported the deviations of the PatchTST in the full-shot setting as a comparison. It can be observed that ROSE exhibits stable performance.
Table 11: Results deviation.
Models ROSE ROSE (10%) PatchTST confidence interval
Metric MSE MAE MSE MAE MSE MAE -
ETTm1 0.342$\pm$0.003 0.367$\pm$0.002 0.349$\pm$0.003 0.372$\pm$0.002 0.349$\pm$ 0.004 0.383$\pm$0.003 99%
ETTm2 0.246$\pm$0.002 0.303$\pm$0.004 0.249$\pm$0.002 0.308$\pm$0.002 0.255$\pm$0.002 0.314$\pm$0.003 99%
ETTh1 0.392$\pm$0.004 0.413$\pm$0.004 0.397$\pm$0.003 0.419$\pm$0.003 0.411$\pm$0.003 0.432$\pm$0.005 99%
ETTh2 0.330$\pm$0.003 0.374$\pm$0.002 0.335$\pm$0.004 0.380$\pm$0.003 0.348$\pm$0.004 0.390$\pm$0.004 99%
Traffic 0.391$\pm$0.008 0.266$\pm$0.005 0.418$\pm$0.011 0.278$\pm$0.006 0.404$\pm$0.009 0.283$\pm$0.002 99%
Weather 0.217$\pm$0.008 0.250$\pm$0.007 0.224$\pm$0.007 0.252$\pm$0.009 0.223$\pm$0.011 0.263$\pm$0.014 99%
Electricity 0.156$\pm$0.007 0.249$\pm$0.009 0.164$\pm$0.004 0.253$\pm$0.004 0.163$\pm$0.009 0.261$\pm$0.013 99%
A.9 Visualization
A.9.1 Visualization analysis
To showcase the benefits of cross-domain pre-training, we performed visualizations in both the zero-shot setting and full-shot setting.
Zero-shot: We pre-train the baselines iTransformer and PatchTST on the energy domain dataset ETTm1 and test their zero-shot performance on two different domains (weather, traffic) . ROSE, without fine-tuning, is evaluated to the same two test-sets. As shown in the Figure 10, we find that the baselines generally perform worse during domain shift due to their poor generalization. However, ROSE excels in scenarios across all domains, which demonstrates the benefits of cross-domain pre-training for improving generalization.
Full-shot: We train the baselines on the train-set of downstream dataset ETTh2 and fine-tune ROSE on the same train-set. As shown in the Figure 11, We find that the baselines is limited by data diversity, leading to poor performance on patterns which rarely appear. However, ROSE excels in these cases, as the cross-domain pre-training allows ROSE to learn diverse temporal patterns, and helps ROSE to predict the patterns which rarely appear in the downstream train-set well.
Figure 10: Visualization comparison of ROSE with cross-domain pre-training and other SOTA baselines in the zero-shot setting for three domain datasets.
Figure 11: Visualization comparison of ROSE with cross-domain pre-training and other SOTA baselines in the full-shot setting for rare and common patterns.
A.9.2 Visualization showcase
To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 12 to Figures 15, given by the following models: DLinear [Zeng et al., 2023], TimesNet [Wu et al., 2022], iTransfomrer [Liu et al., 2023], and PatchTST [Nie et al., 2022]. Among the methods, ROSE demonstrates the most accurate prediction ability.
Figure 12: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset in full-shot setting.
Figure 13: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset in 10% few-shot setting.
Figure 14: Visualization of input-512 and predict-336 forecasting results on the weather dataset in full-shot setting.
Figure 15: Visualization of input-512 and predict-336 forecasting results on the weather dataset in 10% few-shot setting.
A.10 Full Results
A.10.1 Full-shot results
@@
Models ROSE ITransformer PatchTST Timesnet Dlinear GPT4TS S2P-LLM
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.354 0.385 0.386 0.405 0.370 0.400 0.470 0.470 0.367 0.396 0.376 0.397 0.366 0.396
192 0.389 0.407 0.424 0.440 0.413 0.429 0.568 0.523 0.400 0.417 0.416 0.418 0.401 0.420
336 0.406 0.422 0.449 0.460 0.422 0.440 0.595 0.547 0.428 0.439 0.442 0.433 0.412 0.431
720 0.413 0.443 0.495 0.487 0.447 0.468 0.694 0.591 0.468 0.491 0.477 0.456 0.440 0.458
avg 0.391 0.414 0.439 0.448 0.413 0.434 0.582 0.533 0.416 0.436 0.427 0.426 0.406 0.427
ETTh2 96 0.265 0.320 0.297 0.348 0.274 0.337 0.351 0.399 0.302 0.368 0.285 0.342 0.278 0.340
192 0.328 0.369 0.371 0.403 0.341 0.382 0.394 0.429 0.404 0.433 0.354 0.389 0.246 0.385
336 0.353 0.391 0.404 0.428 0.329 0.384 0.415 0.443 0.511 0.498 0.373 0.407 0.367 0.406
720 0.376 0.417 0.424 0.444 0.379 0.422 0.477 0.481 0.815 0.640 0.406 0.441 0.400 0.436
avg 0.331 0.374 0.374 0.406 0.331 0.381 0.409 0.438 0.508 0.485 0.354 0.394 0.347 0.391
ETTm1 96 0.275 0.328 0.300 0.353 0.293 0.346 0.405 0.421 0.303 0.346 0.292 0.262 0.288 0.346
192 0.324 0.358 0.345 0.382 0.333 0.370 0.508 0.473 0.335 0.365 0.332 0.301 0.323 0.365
336 0.354 0.377 0.374 0.398 0.369 0.392 0.523 0.479 0.365 0.384 0.366 0.341 0.359 0.390
720 0.411 0.407 0.429 0.430 0.416 0.420 0.523 0.484 0.418 0.415 0.417 0.401 0.403 0.418
avg 0.341 0.367 0.362 0.391 0.353 0.382 0.490 0.464 0.356 0.378 0.352 0.383 0.343 0.379
ETTm2 96 0.157 0.243 0.175 0.266 0.166 0.256 0.233 0.305 0.164 0.255 0.173 0.262 0.165 0.257
192 0.213 0.283 0.242 0.312 0.223 0.296 0.265 0.328 0.224 0.304 0.229 0.301 0.222 0.299
336 0.266 0.319 0.282 0.340 0.274 0.329 0.379 0.392 0.277 0.339 0.286 0.341 0.277 0.330
720 0.347 0.373 0.378 0.398 0.362 0.385 0.390 0.407 0.371 0.401 0.378 0.401 0.363 0.390
avg 0.246 0.305 0.269 0.329 0.256 0.317 0.317 0.358 0.259 0.325 0.266 0.326 0.257 0.319
Weather 96 0.145 0.182 0.159 0.208 0.149 0.198 0.193 0.244 0.170 0.230 0.162 0.212 0.145 0.195
192 0.183 0.226 0.200 0.248 0.194 0.241 0.320 0.329 0.212 0.267 0.204 0.248 0.190 0.235
336 0.232 0.267 0.253 0.289 0.245 0.282 0.363 0.366 0.257 0.305 0.254 0.286 0.243 0.280
720 0.309 0.327 0.321 0.338 0.314 0.334 0.440 0.404 0.318 0.356 0.326 0.337 0.312 0.326
avg 0.217 0.251 0.233 0.271 0.226 0.264 0.329 0.336 0.239 0.289 0.237 0.270 0.222 0.259
Electricity 96 0.125 0.220 0.138 0.237 0.129 0.222 0.182 0.287 0.141 0.241 0.139 0.238 0.135 0.230
192 0.142 0.235 0.157 0.256 0.147 0.240 0.193 0.293 0.154 0.254 0.153 0.251 0.149 0.247
336 0.162 0.252 0.167 0.264 0.163 0.259 0.196 0.298 0.168 0.271 0.169 0.266 0.167 0.266
720 0.191 0.284 0.194 0.286 0.197 0.290 0.209 0.307 0.203 0.303 0.206 0.297 0.200 0.287
avg 0.155 0.248 0.164 0.261 0.159 0.253 0.195 0.296 0.166 0.267 0.167 0.263 0.161 0.257
Traffic 96 0.354 0.252 0.363 0.265 0.360 0.249 0.611 0.323 0.411 0.294 0.388 0.282 0.379 0.274
192 0.377 0.257 0.385 0.273 0.379 0.256 0.609 0.327 0.421 0.298 0.407 0.290 0.397 0.282
336 0.396 0.262 0.396 0.277 0.392 0.264 0.616 0.335 0.431 0.304 0.412 0.294 0.407 0.289
720 0.434 0.283 0.445 0.312 0.432 0.286 0.656 0.349 0.468 0.325 0.450 0.312 0.440 0.301
avg 0.390 0.264 0.397 0.282 0.391 0.264 0.623 0.333 0.433 0.305 0.414 0.294 0.405 0.286
A.10.2 Few-shot results
@@
Models ROSE ITransformer PatchTST Timesnet Dlinear GPT4TS S2P-LLM
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.367 0.395 0.442 0.464 0.458 0.463 0.579 0.522 1.355 0.816 0.458 0.456 0.481 0.474
192 0.399 0.416 0.476 0.475 0.481 0.490 0.641 0.553 1.210 0.825 0.570 0.516 0.518 0.491
336 0.405 0.423 0.486 0.482 0.465 0.475 0.721 0.582 1.487 0.914 0.608 0.535 0.664 0.570
720 0.416 0.443 0.509 0.506 0.478 0.492 0.630 0.574 1.369 0.826 0.725 0.591 0.711 0.584
avg 0.397 0.419 0.478 0.482 0.470 0.480 0.643 0.558 1.355 0.845 0.590 0.525 0.593 0.529
ETTh2 96 0.273 0.332 0.333 0.385 0.350 0.389 0.378 0.413 1.628 0.724 0.331 0.374 0.354 0.400
192 0.334 0.376 0.402 0.428 0.416 0.426 0.463 0.460 1.388 0.713 0.402 0.411 0.400 0.423
336 0.358 0.397 0.438 0.452 0.401 0.429 0.507 0.495 1.595 0.772 0.406 0.433 0.442 0.450
720 0.376 0.417 0.466 0.477 0.436 0.457 0.516 0.501 1.664 0.857 0.449 0.464 0.480 0.486
avg 0.335 0.380 0.410 0.436 0.401 0.425 0.466 0.467 1.569 0.766 0.397 0.421 0.419 0.439
ETTm1 96 0.287 0.336 0.353 0.392 0.317 0.363 0.481 0.446 0.454 0.475 0.390 0.404 0.388 0.401
192 0.331 0.362 0.385 0.410 0.351 0.382 0.621 0.491 0.575 0.548 0.429 0.423 0.422 0.421
336 0.362 0.379 0.422 0.432 0.376 0.398 0.521 0.479 0.773 0.631 0.469 0.439 0.456 0.430
720 0.416 0.412 0.494 0.472 0.435 0.430 0.571 0.508 0.943 0.716 0.569 0.498 0.554 0.490
avg 0.349 0.372 0.414 0.426 0.370 0.393 0.549 0.481 0.686 0.593 0.464 0.441 0.455 0.435
ETTm2 96 0.159 0.247 0.183 0.279 0.170 0.259 0.212 0.292 0.493 0.476 0.188 0.269 0.192 0.274
192 0.217 0.287 0.247 0.320 0.226 0.297 0.297 0.353 0.923 0.658 0.251 0.309 0.246 0.313
336 0.269 0.322 0.300 0.353 0.284 0.333 0.328 0.364 1.407 0.822 0.307 0.346 0.301 0.340
720 0.357 0.377 0.385 0.408 0.363 0.382 0.456 0.440 1.626 0.905 0.426 0.417 0.400 0.403
avg 0.250 0.308 0.279 0.340 0.261 0.318 0.323 0.362 1.112 0.715 0.293 0.335 0.284 0.332
Weather 96 0.145 0.184 0.189 0.229 0.166 0.217 0.199 0.248 0.230 0.318 0.163 0.215 0.159 0.210
192 0.190 0.227 0.239 0.269 0.211 0.257 0.249 0.285 0.357 0.425 0.210 0.254 0.200 0.251
336 0.245 0.269 0.294 0.308 0.261 0.296 0.297 0.316 0.464 0.493 0.256 0.292 0.257 0.293
720 0.317 0.328 0.366 0.356 0.328 0.342 0.367 0.361 0.515 0.532 0.321 0.339 0.317 0.335
avg 0.224 0.252 0.272 0.291 0.242 0.278 0.278 0.303 0.391 0.442 0.238 0.275 0.233 0.272
Electricity 96 0.135 0.226 0.184 0.276 0.161 0.256 0.279 0.359 0.227 0.334 0.139 0.237 0.143 0.243
192 0.150 0.240 0.192 0.284 0.163 0.257 0.282 0.363 0.265 0.366 0.156 0.252 0.159 0.258
336 0.166 0.258 0.216 0.308 0.173 0.266 0.289 0.367 0.339 0.417 0.175 0.270 0.170 0.269
720 0.205 0.290 0.265 0.347 0.221 0.313 0.333 0.399 0.482 0.478 0.233 0.317 0.230 0.315
avg 0.164 0.253 0.214 0.304 0.180 0.273 0.296 0.372 0.328 0.399 0.176 0.269 0.175 0.271
Traffic 96 0.398 0.270 0.458 0.314 0.421 0.299 0.705 0.386 0.616 0.385 0.414 0.297 0.403 0.293
192 0.405 0.270 0.473 0.319 0.439 0.313 0.710 0.393 0.710 0.480 0.426 0.301 0.412 0.295
336 0.417 0.277 0.491 0.329 0.448 0.318 0.863 0.456 0.723 0.481 0.434 0.303 0.427 0.316
720 0.452 0.294 0.536 0.361 0.478 0.320 0.928 0.485 0.673 0.436 0.487 0.337 0.469 0.325
avg 0.418 0.278 0.490 0.331 0.447 0.312 0.801 0.430 0.680 0.446 0.440 0.310 0.427 0.307
A.10.3 Ablation study results
Novelty of decomposed frequency learning. Frequency masking is not a new concept, but past approaches randomly mask frequencies of a single time series once  [Chen et al., 2023b, Zhang et al., 2023], which show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains. While the multi-frequency masking we proposed randomly mask either high-frequency or low-frequency components of a time series multiple times as the key to enable learning of common time series patterns, such as trends and various long and short term fluctuations. Moreover, different from utilizing frequency masking as a way of data augmentation to enhance the diversity of input data  [Chen et al., 2023b, Zhang et al., 2023], we combine multi-frequency masking with reconstruction task as a novel pre-training framework, that learns a universal and unified feature representation by comprehending the data from various frequency perspectives, thereby enabling it to learn generalized representations.
Difference between frequency-domain masking and time-domain noise addition. Multi-frequency masking and reconstruction are not equivalent to the pre-training methods of adding noise and denoising (Noise). Due to the sparsity of time series, the process of adding noise and denoising may potentially disrupt the information of original time series [Dong et al., 2024]. In contrast, multi-frequency masking not only preserves the series from such disruption but also helps the model understand temporal patterns from a multi-frequency perspective, thereby helping the model to learn general features better.
Other pre-training tasks. Based on the two points above, we conduct experiments to compare two other pre-training tasks: 1) using frequency-domain augmentation only for data expansion without reconstruction task (Aug); 2) replacing multi-frequency masking and reconstruction task with adding time-domian noise and denoise task (Noise).As shown in Table 14, we find that ROSE is significantly more effective than Aug and Noise, which demonstrates the effectiveness of multi-frequency masking and reconstruction task in learning generalized features.